OpenStudy (rayman5766):

Simplify. –12 – 4.8 ∙ (–2 + –8) A. –168 B. 36 C. 72 D. 168

OpenStudy (anonymous):

- 12 - 4.8 * (- 2 + - 8) = - 12 - (4.8 * (-10) = -12 - (- 48) = -12 + 48 = B. 36